OpenStudy (anonymous):

Write the rule for the following arithmetic sequence: 17, 22, 27, 32, ... a] tn=5n+16 b] tn=20n+480 c] tn=-5n+12 d] tn=5n+31 i don't understand what to do

OpenStudy (anonymous):

Let see: \(\Large t_0= 17=17+0\times5\ \ (n=0)\\\Large t_1=22=17+5\ \ (n=1)\\\Large t_2=27=22+5=17+2\times5\ \ (n=2)\\\Large t_3=32=27+5=17+3\times5\ \ (n=3)\) Thus, the answer is \(\Large tn=5n+17\)